OpenStudy (anonymous):

a set of data has a normal distribution with a mean of 120 and a standard deviation of 10. Find the interval about the mean within 90% of the data lie

OpenStudy (amistre64):

since the mean is 50% of the data, and 90/2 = 45 you want z values that correspond to 50-45 and 50+45

OpenStudy (amistre64):

i take it that made perfect sense :)

OpenStudy (amistre64):

z1 = -1.645 z2 = 1.645 \[z_n=\frac{x-mean}{sd}\] \[z_n(sd)+mean=x\] \[\pm1.645(10+120)=x_1,x_2\]

OpenStudy (amistre64):

latex fail :) \[\pm1.645(10)+120=x_1,x_2\]

OpenStudy (anonymous):

so the choices are: 94.2-145.8 103.5-136.5 113.68-126.32 100.4-139.6

OpenStudy (amistre64):

so about, 120 - 16 and 120 + 16

OpenStudy (anonymous):

ohh thanks!